OpenStudy (anonymous):

3(1-2k)=-5k Help please like help step by step.

OpenStudy (nikato):

Distribute that 3 first

OpenStudy (anonymous):

??

OpenStudy (australopithecus):

Distribution Rule: \[A(B + C) = AB + AC\]

OpenStudy (the_fizicx99):

3(1-2k)=-5k 3*1 and 3*(-2k) 3 -6k = -5k +6k +6k 3 = 1k or 3 = k

OpenStudy (nikato):

Distribute like this a(b-c)= ab-ac

OpenStudy (anonymous):

Oh gosh so many people at one @nikato help!

OpenStudy (nikato):

Ok, can u distribute the left side?

OpenStudy (anonymous):

3x 1 right and that's 3

OpenStudy (nikato):

Yes

OpenStudy (nikato):

And what about the other number in the parantheses

OpenStudy (anonymous):

Would you subtract 3 - 2?

OpenStudy (nikato):

No, so distribute. A(B+C)=AB+AC so now using ur numbers 3(1-2k)= 3(1) - 3(2k)so that is 3-6k Do u understand this part now?

OpenStudy (anonymous):

Oh so now we add 3 +2??? Right I proubley annoy you

OpenStudy (nikato):

No, u would not ever have to add 3+2

OpenStudy (nikato):

So after distributing u will have

OpenStudy (nikato):

3-6k=-5k Can u solve for k from here now?

OpenStudy (anonymous):

yeah ...

OpenStudy (nikato):

Ok, and of u did it correctly, k=3